OpenStudy (anonymous):

plz help will fan and medal. Jared has a wooden board that is 11 ft long. He wants to cut it into two pieces so that the longer piece is 4 ft less than 4 times the length of the shorter piece. How long will each piece of board be?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Wiild_pickle

OpenStudy (anonymous):

plz help asap

OpenStudy (luffingsails):

Try setting up the equation. Let x = shorter piece The other piece is y = 4 * x - 4 (4 feet shorter than 4 times the short piece) The two pieces together total 11 ft. So, x + 4 * x - 4 = 11 5x = 7 x = 7/5 = 1.4 ft. y = 4 * 7/5 - 4 = 1.6 ft

OpenStudy (anonymous):

thanks

OpenStudy (luffingsails):

You're welcome!

OpenStudy (anonymous):

The equation is correct, but the answer is not correct. The second step has an error.